3x+y+2z=80 x+2y-z=10 2x-y+3z=70
solve please

Answer by stanbon(75887)

3x+y+2z=80
x+2y-z=10
2x-y+3z=70
----
Use any method you know to get:
x = 30
y = -10
z = 0
==========
Cheers,
Stan H.